QUESTION 6 Holding all else constant, as additional firms enter an industry O a.more output is available at each given price O b. less output is available at each given price O c. the same output is available at each given price O d. Unable to tell QUESTION 7 Suppose market demand andmarket supply are given by Qd 15-4P and Qs What are the equilibrium quantity and price in this market? Show your work? -3+2P
Answer #1

Solution -

Question 6

Option A.more output is available at each given price. is Correct Answer.

Reason -

when the holding at that time output will stop when next firm enters in market at that time more output will available at each price so above given option is correct option .

Question 7

we know
Equlibrium Condtion

Qd = Qs

15 - 4P = -3+2P

6P = 18

P = $3

Now Put P = $3 in Qd For Equilibrium Quantity

15 - 4P = Q

Q =15 - 12

Q = 3

NOTE - AS PER HOMEWORKLIB POLICY WE CAN SOLVE TOP 1 QESTION BUT HERE I SOLVE BOTH FOR YOU

DO UP VOTE HAVE A NICE DAY .